# MoSE: Hierarchical Self-Distillation Enhances Early Layer Embeddings

### v2

@@ Abstract
Deploying language models often requires navigating accuracy vs. performance trade-offs to meet latency constraints while preserving utility. Traditional model distillation reduces size but incurs substantial costs through training separate models. We introduce ModularStarEncoder (MoSE), a 1-billion-parameter multi-exit encoder for code retrieval and classification that employs a novel Self-Distillation mechanism. This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs. Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training—improving intermediate representations at minimal additional cost. We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization. Additionally, we release a new dataset created through code translation that extends text-to-code benchmarks with cross-language code-to-code pairs. Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
@@ 1 Introduction
Large language models (LLMs) have significantly impacted natural language processing (Niu et al., 2023), but their computational demands pose substantial deployment challenges. The field has responded with various efficiency strategies: quantization techniques reduce numerical precision (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al., 2019; Jiao et al., 2019); and pruning methods eliminate less influential weights (Han et al., 2015). Concurrently, model families like LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks. Early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al., 2022) enables adjustment of computational complexity through embedding dimensionality pruning. These approaches maintain performance while reducing computational demands in resource-constrained environments. Furthermore, insights from the SimCLR framework (Chen et al., 2020) revealed that the final layer of a model does not necessarily produce the most semantically meaningful representations. Valeriani et al. (2023) analyzed the geometric evolution of hidden representations within large transformer models across layers, outlining that intermediate layers are identified as holding the most semantically rich representations for downstream tasks.
@@
Figure 1: (a) Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$. (b) Computational cost (GFLOPs) vs. performances trade-off at different exit layers over CodeSearchNet dataset (text-to-code in avg. MRR) and POJ104 (code-to-code in mAP). Despite a reduction of approximately 90% floating point operations from layer 36 to layer 4, MRR performance only drops by 6.4% in absolute terms. For POJ104, our best results are observed in the initial layers.
Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al., 2024) architecture, with 1 billion parameters that incorporate intra-model self-distillation. As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective. This approach encourages lower layers to learn better representations by mimicking higher-layer outputs. Figure 1(b) demonstrates the resulting accuracy-computation trade-off of encoders trained under this framework: despite a 90% reduction in floating point operations from layer 36 to layer 4, performance over text-to-code retrieval drops by only 6.4% in absolute terms, while the model maintains the best performances for code-to-code retrieval in very early layers.
Notably, the exit point with the best performance isn’t necessarily the final layer, as noted in previous findings due to Chen et al. (2020); Valeriani et al. (2023). This offers both deployment flexibility and a natural model selection mechanism for specific tasks. Users can select which portion of the model to use during inference, starting from the first exit point at 160M parameters. By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands. This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al., 2023).
We fine-tuned MoSE on SynthCoNL (cf. Sec. 5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko & Roy, 2015) for code clone detection, achieving state-of-the-art results among open models while maintaining deployment modularity.
@@
    A self-distillation methodology that trains multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability—an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
@@
    MoSE, pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
@@
    SynthCoNL [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated. Positive examples originate from the same repository context, whereas negative examples are sourced from different repositories. To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence. On the right side, you’ll find the pseudocode for the in-context loss framework.
@@ 2 Self-Distillation and In-Context Classification
As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al., 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability. Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
We hypothesize that applying supervision at multiple depths, not just at the output—encourages more robust and semantically rich intermediate representations. To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al., 2023). While we retain the widely adopted Masked Language Modeling (MLM) loss Feng et al. (2020), we discard the traditional Next Sentence Prediction (NSP) objective. NSP has been shown to offer minimal benefit after fine-tuning (Liu et al., 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially for long-form code inputs. In its place, we introduce an In-Context Classification (ICC) loss (see section 2.1) that increases input density by reducing padding and fosters better chunk-level comprehension, essential for tasks like semantic retrieval or classification.
@@ 2.1 Masked Language Modeling and In-Context Classification
[…]
Table 1: Performance—reported in MRR, and when specified in NDCG—of different models on text-to-code with CodeSearchNet using CodeXGLUE. We reported the results presented in CodeT5+, UniXcoder and ModernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024). MoSE surpasses all existing open-source models, achieving an improvement of +3.6% compared to the best-performing open model, CodeT5+. MoSE reduced the performance gap between open-source models and closed-source ones such as OpenAI text-embedding-3-large.
CodeSearchNet
Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
CodeT5+ (770M) 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
ModernBERT-large - - - - - - - 59.5
OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
@@ 2.2 Self-Distillation: Multi-Layer loss
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This Self-Distillation mechanism, shown in Figure 1, allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
@@ 3 Pre-training
We built MoSE on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its original size, obtaining a model of 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. We have included and summarized additional details about model architecture in appendix D.
@@ 3.1 Training Details
We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al., 2024) dataset.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$. We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al., 2024) with 4,000 warmup steps. The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU working hours on the Leonardo supercomputer (Turisini et al., 2023).
@@ 4 Downstream tasks
4.1 Finetuning for Retrieval, text-to-code and code-to-code search
Using the SynthCoNL dataset detailed in section 5, we leveraged a instruction prompting approach Su et al. (2023) and trained a single model for both text-to-code and code-to-code retrieval tasks. The optimization objective uses CLIP loss (Radford et al., 2021) with the same multi-layer approach discussed in section 2.2, modified to enhance representation learning. Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2,048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with the state-of-the-art.
@@ 4.1 Finetuning for Retrieval, text-to-code and code-to-code search
4.2 Finetuning for classification, code clone detection
[…]

### v3

@@ Abstract
Deploying language models often requires navigating accuracy vs. performance trade-offs to meet latency constraints while preserving utility. Traditional model distillation reduces size but incurs substantial costs through training separate models. We introduce ModularStarEncoder (MoSE), a 1-billion-parameter multi-exit encoder for code retrieval and classification that employs a novel Self-Distillation mechanism. This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs. Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training, thereby improving intermediate representations at minimal additional cost. We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization. Additionally, we release a new dataset created through code translation that extends text-to-code benchmarks with cross-language code-to-code pairs. Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
MoSEhttps://huggingface.co/modularStarEncoder
SynthCoNLhttps://huggingface.co/datasets/modularStarEncoder/SynthCoNL
@@ 1 Introduction
Large language models (LLMs) have significantly impacted natural language processing (Niu et al. 2023), but their computational demands pose substantial deployment challenges. The field has responded with various strategies aimed at improving efficiency: quantization techniques reduce numerical precision (Jacob et al. 2017; Lin et al. 2023; Egiazarian et al. 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al. 2019; Jiao et al. 2019); and pruning methods eliminate less influential parameters (Han et al. 2015). Concurrently, model families like LLaMA (Dubey et al. 2024), Qwen (Hui et al. 2024), Mistral (Jiang et al. 2023), and SmolLM (Allal et al. 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks. Early-exit architectures such as BranchyNet (Teerapittayanon et al. 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al. 2022) enables adjustment of computational complexity through embedding dimensionality pruning. These approaches maintain performance while reducing computational demands in resource-constrained environments. Furthermore, insights from the SimCLR framework (Chen et al. 2020) revealed that the final layer of a model does not necessarily produce the most useful representations. Valeriani et al. (2023) analyzed the geometric evolution of hidden representations within large transformer models across layers, outlining that intermediate layers are identified as holding the most semantically rich representations for downstream tasks.
@@
Figure 1: (a) Overview of our multi‐exit Self-Distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$. (b) Computational cost (GFLOPs) vs. performances trade-off at different exit layers over CodeSearchNet dataset (text-to-code in avg. MRR) and POJ104 (code-to-code in mAP). Despite a reduction of approximately 90% floating point operations from layer 36 to layer 4, MRR performance only drops by 6.4% in absolute terms. For POJ104, our best results are observed in the initial layers.
Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al. 2024) architecture, with 1 billion parameters that incorporate intra-model Self-Distillation. As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective. This approach encourages lower layers to learn better representations by mimicking higher-layer outputs. Figure 1(b) demonstrates the resulting accuracy-computation trade-off of encoders trained under this framework: despite a 90% reduction in floating point operations from layer 36 to layer 4, performance over text-to-code retrieval drops by only 6.4% in absolute terms, while the model maintains the best performances for code-to-code retrieval in very early layers.
Notably, the exit point with the best performance isn’t necessarily the final layer, as noted in previous findings due to Chen et al. (2020); Valeriani et al. (2023). This offers both deployment flexibility and a natural model selection mechanism for specific tasks. Users can select which portion of the model to use during inference, starting from the first exit point at 160M parameters. By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands. This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al. 2023).
We fine-tuned MoSE on SynthCoNL (cf. Sec. 5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko and Roy 2015) for code clone detection, achieving competitive results among open models while maintaining deployment modularity.
@@
    A Self-Distillation methodology that trains multiple models within a unified layer stack, reducing redundancy and improving scalability, an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
@@
    MoSE, pre-trained and fine-tuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
@@
    SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated. Positive examples originate from the same repository context, whereas negative examples are sourced from different repositories. To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence. On the right side, you’$l$ find the pseudocode for the in-context loss framework.
@@ 2 Self-Distillation and In-Context Classification
As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al. 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability. Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
We hypothesize that applying supervision at multiple depths, not just at the output, encourages more robust and semantically rich intermediate representations. To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al. 2023). While we retain the widely adopted Masked Language Modeling (MLM) loss (Feng et al. 2020a), we discard the traditional Next Sentence Prediction (NSP) objective. NSP has been shown to offer minimal benefit after fine-tuning (Liu et al. 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially so for long-form code inputs. In its place, we introduce an In-Context Classification (ICC) loss (see Section 2.1) that increases input density, reducing padding and fostering better chunk-level comprehension, essential for downstream tasks such as semantic retrieval or classification.
@@ 2.1 Masked Language Modeling and In-Context Classification
In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head. Following Zhang et al. (2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy (Devlin et al. 2019). The secondary objective, $\mathcal{L}_{ICC}$, classifies whether randomly concatenated inputs (separated by a $\text{SEP}]$ special token) originate from the same repository (see Figure 2). Each concatenated sample has a 50% probability of containing source code from different repositories. This approach increases input density, reducing padding by expanding the average input length from 630 to 1,300 tokens. Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may also improve inter-language understanding.
CodeSearchNet CT POJ104
Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG MRR mAP
MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2 98.9 75.9
CodeT5+ 78.0 71.3 92.7 75.8 76.2 70.1 77.4 - 98.4 24.5
UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 - 97.6 41.0
ModernBERT-large - - - - - - - 59.5 93.1 27.3
OpenAI Embedding 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3 98.8 82.9
Table 1: Performance of different models on text-to-code and code-to-code tasks using the CodeXGLUE benchmark suite, specifically in terms of MRR (on CodeSearchNet and CodeTranslation CT dataset), NDCG, and mAP (on POJ104). On CodeSearchNet, MoSE surpasses all existing open-source models, achieving an improvement of 3.6% compared to the best-performing open model, CodeT5+. Furthermore, MoSE significantly reduces the performance gap between open-source models and closed-source alternatives, such as OpenAI text-embedding-3-large referred to as OpenAI Embedding, and performs comparably on the CT dataset. MoSE demonstrates competitive results with the CT benchmark and also outperforms all open-source models on the POJ104 benchmark.
@@ 2.2 Self-Distillation: Multi-Layer loss
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (Section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha\$where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36). We also train a “tiny” single-exit baseline to enable isolating the Self-Distillation effect. The Self-Distillation mechanism illustrated in Figure 1 allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
@@ 3 Pre-training
We built MoSE on top of StarCoder-2 (Lozhkov et al. 2024), applying several modifications to the model. We reduced its original size, resulting in a 1B-parameter model. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al. 2023) with 16 attention heads and 4 key-value heads. MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al. 2021) with a base period $\theta=10^{6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al. 2021). Finally, our implementation integrates FlashAttention V2 (Dao 2023) for faster inference. We have included and summarized additional details about the model architecture in the Appendix.
@@ 3.1 Training Details
We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al. 2024) dataset.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$. We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al. 2024) with 4,000 warmup steps. The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU hours.
@@ 4 Downstream Tasks
4.1 Fine-tuning for Retrieval, text-to-code and code-to-code search
Using the SynthCoNL dataset detailed in Section 5, we leveraged an instruction prompting approach (Su et al. 2023) and trained a single model for both text-to-code and code-to-code retrieval tasks. The optimization objective uses CLIP-style loss (Radford et al. 2021) with the same multi-layer approach discussed in Section 2.2, modified to enhance representation learning. Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2,048 examples, ensuring that text-to-code and code-to-code were equally distributed across the batch.
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the fine-tuning process, and set the temperature parameter at 10.0.
@@ 4.1 Fine-tuning for Retrieval, text-to-code and code-to-code search
Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with open models.
4.2 Fine-tuning for classification, code clone detection
In addition to fine-tuning the model for multiple retrieval tasks as discussed in Section 4.1, we also explored how our methodology performs in a classification setup. We tested how the Multi-Layer and the In-Context classification losses adapt to the code clone detection task by: (1) fine-tuning multiple classification heads, one for each of the five exit points (layers 4, 9, 18, 27, and 36), and (2) leveraging the model’s ability to understand relationships between code snippets by inputting pairs of (hypothetical clone) code segments and training it to classify whether they are clones. We followed the input pattern devised in Section 2.1, formatting the input as follows: $\text{SEP}]$ snippet-1 $\text{SEP}]$ snippet-2 $\text{CLS}]$, and then used the final $\text{CLS}]$ token representation as input to a classifier. We fine-tuned the model with a constant learning rate of $1\mathrm{e}{-5}$ with 2,000 warmup steps, and used a batch size of 64 elements and trained for 14,000 steps.